40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 191 of 354  Not logged in ELOG logo
ID Date Author Type Category Subject
  8244   Wed Mar 6 18:51:07 2013 AnnalisaUpdateAlignmentAuxiliary laser installed for FSR and TMS measurement of the PRC

We want to measure the g-factor of the PRC using the beat note of the main laser with an auxiliary NPRO laser.

We are going to phase lock the NPRO to the main laser (taking it from POY) and then we will inject the NPRO  through the AS edge of the ITMY.

Today Sendhil and I installed the auxiliary laser on the ITMY table moving it from the AS table.

We also installed the beam steering optics, except the BS which will launch the beam through the AR edge of the ITMY.

To do: install the BS, take the POY beam and mix it with the auxiliary laser on a photodiode to phase lock the two lasers, do better calculations for the mode matching optics to be used for the auxiliary laser beam.

Attachment 1: IMG_3-6-13.JPG
IMG_3-6-13.JPG
  8243   Wed Mar 6 18:27:24 2013 JamieUpdateGeneralnow recording input TT channels to frames, but why no autoburt?

I spent some time trying to figure out how to get a record of the pointing of the input pointing tip-tilt (TT) channels.

Frames

Currently the TT pointing is done via the offset in the PIT/YAW filter banks, ie. C1:IOO-TT1_PIT_OFFSET, which is an EPICS record.  I added these channels to the C0EDCU.ini, which (I'm pretty sure) specifies which EPICS channels are recorded to frames.

controls@pianosa:~ 0$ grep C1:IOO-TT /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini 
[C1:IOO-TT1_PIT_OFFSET]
[C1:IOO-TT1_YAW_OFFSET]
[C1:IOO-TT2_PIT_OFFSET]
[C1:IOO-TT2_YAW_OFFSET]
controls@pianosa:~ 0$ 

I then confirmed that the data is being recorded:

controls@pianosa:~ 0$ FrChannels /frames/full/10466/C-R-1046657424-16.gwf | grep TT
C1:IOO-TT1_PIT_OFFSET 16
C1:IOO-TT1_YAW_OFFSET 16
C1:IOO-TT2_PIT_OFFSET 16
C1:IOO-TT2_YAW_OFFSET 16
controls@pianosa:~ 0$ 

BURT

The EPICS records for these channels *should* be recorded by autoburt, but Yuta noticed they were not:

controls@pianosa:~ 0$ grep -R C1:IOO-TT1_PIT_OFFSET /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2013/Mar/6/
controls@pianosa:~ 1$

The autoburt log seems to indicate some sort of connection problem:

controls@pianosa:! 130$ grep C1:IOO-TT1_PIT_OFFSET /opt/rtcds/caltech/c1/burt/autoburt/logs/c1assepics.log
pv >C1:IOO-TT1_PIT_OFFSET< nreq=-1
pv >C1:IOO-TT1_PIT_OFFSET.HSV< nreq=-1
pv >C1:IOO-TT1_PIT_OFFSET.LSV< nreq=-1
pv >C1:IOO-TT1_PIT_OFFSET.HIGH< nreq=-1
pv >C1:IOO-TT1_PIT_OFFSET.LOW< nreq=-1
C1:IOO-TT1_PIT_OFFSET ... ca_search_and_connect() ... OK
C1:IOO-TT1_PIT_OFFSET.HSV ... ca_search_and_connect() ... OK
C1:IOO-TT1_PIT_OFFSET.LSV ... ca_search_and_connect() ... OK
C1:IOO-TT1_PIT_OFFSET.HIGH ... ca_search_and_connect() ... OK
C1:IOO-TT1_PIT_OFFSET.LOW ... ca_search_and_connect() ... OK
C1:IOO-TT1_PIT_OFFSET ... not connected so no ca_array_get_callback()
C1:IOO-TT1_PIT_OFFSET.HSV ... not connected so no ca_array_get_callback()
C1:IOO-TT1_PIT_OFFSET.LSV ... not connected so no ca_array_get_callback()
C1:IOO-TT1_PIT_OFFSET.HIGH ... not connected so no ca_array_get_callback()
C1:IOO-TT1_PIT_OFFSET.LOW ... not connected so no ca_array_get_callback()
controls@pianosa:~ 0$ 

This is in contrast to successfully recorded channels:

controls@pianosa:~ 0$ grep C1:LSC-DARM_OFFSET /opt/rtcds/caltech/c1/burt/autoburt/logs/c1lscepics.log
pv >C1:LSC-DARM_OFFSET< nreq=-1
pv >C1:LSC-DARM_OFFSET.HSV< nreq=-1
pv >C1:LSC-DARM_OFFSET.LSV< nreq=-1
pv >C1:LSC-DARM_OFFSET.HIGH< nreq=-1
pv >C1:LSC-DARM_OFFSET.LOW< nreq=-1
C1:LSC-DARM_OFFSET ... ca_search_and_connect() ... OK
C1:LSC-DARM_OFFSET.HSV ... ca_search_and_connect() ... OK
C1:LSC-DARM_OFFSET.LSV ... ca_search_and_connect() ... OK
C1:LSC-DARM_OFFSET.HIGH ... ca_search_and_connect() ... OK
C1:LSC-DARM_OFFSET.LOW ... ca_search_and_connect() ... OK
C1:LSC-DARM_OFFSET ... ca_array_get_callback() nreq 1 ... OK
C1:LSC-DARM_OFFSET.HSV ... ca_array_get_callback() nreq 1 ... OK
C1:LSC-DARM_OFFSET.LSV ... ca_array_get_callback() nreq 1 ... OK
C1:LSC-DARM_OFFSET.HIGH ... ca_array_get_callback() nreq 1 ... OK
C1:LSC-DARM_OFFSET.LOW ... ca_array_get_callback() nreq 1 ... OK
controls@pianosa:~ 0$ 

In fact all the records in the c1assepics log are showing the same "not connected so no ca_array_get_callback()" error.  I don't know what the issue is.  I have no problem reading the values from the command line, with e.g. ezcaread.  So I'm perplexed.

If anyone has any idea why the c1ass EPICS records would fail to autoburt, let me know.

  8242   Wed Mar 6 18:14:33 2013 ManasaSummaryLSCCalibration of BS, ITMX and ITMY actuators

[Yuta, Manasa]

Measured actuator response between 50Hz and 200 Hz in (m/counts).

BS     = (20.7 +/- 0.1)    x 10 -9 / f2

ITMX = (4.70 +/- 0.02) 
x 10 -9/ f2

ITMY = (4.66 +/- 0.02)
x 10 -9/ f2

Actuator response differs by 30% for all the 3 mirrors from the previous measurements made by Kiwamu in 2011.

Calibration of BS, ITMX and ITMY actuators

We calibrated the actuators using the same technique as in Kiwamu's elog.

A) Measure MICH error

1. Locked Y-arm and X-arm looking at TRY and TRX.
2. Misaligned ETMs
3. Measured  MICH error using ASDC and AS55_Q err (MICH_OFFSET = 20 - to compensate for offset in AS_Qerr which exists even after resetting LSC offsets)

MICHerrCalibration.png

B) Open loop transfer function for MICH control

1. Measured the transfer function between C1:LSC-MICH_IN1 and C1:LSC-MICH_IN2 by exciting on  C1:LSC-MICH_EXC.
MICH filter modules used for measurements(0:1 , 2000:1, ELP50). ELP50 used so that actuation signals above 50 Hz are not suppressed.

MICHOLTF.png

C) Calibration of BS/ ITMX/ ITMY actuators

1. Measured transfer function between actuation channels on BS/ ITMX/ ITMY and C1:LSC-AS55_Q_ERR.

actcalibITMBS.png
 

  8241   Wed Mar 6 16:14:27 2013 SteveUpdateSAFETYsafety training

Chloe Ling, Max Horton and Annalisa Allocca have received basic 40m specific safety training.

Attachment 1: IMG_500.jpg
IMG_500.jpg
  8240   Wed Mar 6 11:33:09 2013 steveUpdateSAFETYsafety audit 2013

 

                                                 Recommended correction list:

1,  refill- upgrade first aid boxes

2,  maintain 18" ceiling to bookshelf clearance so the ceiling fire sprinklers are not blocked: room 101

3,  label chilled water supply & return valves in IFO room

4,  calibrate bake room hoods annually

5,  update safety sign at fenced storage

 

              40m still to do list:

1,   clean and measure all safety glasses

2,   annual crane inspection is scheduled for 8am March 19, 1013

3,   make PSL encloser shelf earthquake proof

 

Do you see something that is not safe? Add it to this list please.

 

 

Attachment 1: IMG_0010.JPG
IMG_0010.JPG
  8239   Wed Mar 6 09:44:29 2013 KojiUpdateLSCPRMI locking for g-factor measurement

- What about normalizing POPDC to indicate the carrier recycling gain?

- When you align the PMC, confirm FSS SLOW DC is around zero. Some region of the slow thermal actuation makes the laser source emit at multiple frequencies. In the case, the cavity visibility get worse.

- Do you guys think we can determine if the TT is longitudinally quiet enough? Is there any comparison between the simple Michelson and the PRC motion in m/rtHz?

  8238   Wed Mar 6 08:40:03 2013 SteveUpdateVACfirst scan of this pumpdown

 Pumpdown 75 - Maglev - day 12

Precondition: 66 days at atm, installed TIP-TILTs with coils that  replaced PZT-Jena input steering

 

Attachment 1: pd75m12d.png
pd75m12d.png
Attachment 2: pd75mRGA12d.png
pd75mRGA12d.png
  8237   Wed Mar 6 02:46:20 2013 ManasaUpdateLSCPRMI locking for g-factor measurement

 [Yuta, Manasa]

PRMI alignment procedure for carrier locking has been kept the same except that a couple of issues that have persisted are now taken care of.

We were able to keep PRMI locked for over a minute (POPDC measures 2200) .

1. Trigger levels to MICH and PRCL for PRMI locking have been changed
Whenever we enabled LSC controls to lock PRMI, ITMY moves haphazard if PRM is not aligned. This is because of the low trigger levels of POPDC which keeps MICH triggered all the time while we align PRM. Increasing POPDC trigger (Upper level : 1000 and Lower level:20) for both MICH and PRCL solved this problem and resulted in a more stable ITMY. Also this has stabilized locking greatly if the alignment is fair enough.

2. C1:SUS-ITMY_LSC_GAIN reset
Quoting Yuta's elog "  I found C1:SUS-ITMY_LSC_GAIN is somehow set to be 2.895 recently. I think this should be 1.0. Maybe this is why we had actuation imbalance in ITMs(elog #8212)."
This has been reset to 1.0 and it has not affected PRMI locking.


Mystery

1. Filter module (FM1) on PRCL and MICH show significant delay while enabling and disabling.

2. I tried to fix PMC alignment (PMC trans was 0.76). I was not able to get PMC trans more than 0.79.
PMC has been this way since yesterday.

3. MICH is still bright when locked (ASDC_OUT reads 0.08 for dark and 2.0 for bright). We suspect it is because of the AS55_I error offset that persists even after running LSCoffsets script.

4. PRMI shows some dither at 3Hz when locked.
POPDCspec_PRMIlocked.png

  8236   Tue Mar 5 23:37:11 2013 yutaUpdateSUSPRM angular motion spectra

I measured PRM angular motion spectra (in daytime today).
PRM angular motion is ~ 10 urad in RMS when undamped and ~1 urad in RMS when damped.
If PR2/PR3 angular motions are something like this, and their motion are not enhanced when PRC is locked, measured g-factor of PRC looks OK. But considering the error we have, maybe we are not OK yet. We need calculation.

PRMangularmotion.png

  8235   Tue Mar 5 23:00:08 2013 yutaUpdateLSCYarm and PRC g-factor from misalignment measurement

I fitted intra-cavity power dependance on mirror misalignment plot with parabola to get the g-factor.

  Y arm (tangential) g = 0.44 +0.01 -0.01  (measured value before was 0.3765 +/- 0.003 elog #6938)
  PRC (sagittal)       g = 0.97 +0.01 -0.04 (expected value is 0.939 elog #8068)
  PRC (tangential)   g = 0.96 +0.02 -0.05 (expected value is 0.966 elog #8068)

Error bars are just statistical errors from the fitting. Estimated systematic error is ~0.04 (or more).
Here, I assumed PR2/PR3 to be flat to make the calculation simple. I assumed PRC to be curved PRM - flat ITM cavity, and Y arm to be curved ETMY - flat ITMY cavity.

g-factor calculation:
  Intra-cavity power decrease can be written as

dP/P = (dx/w0)**2 + (dt/a0)**2

where dx and dt are translation and tilt of the beam axis introduced by mirror misalignment. w0 is waist size and a0 is divergence angle (= lamb/(pi*w0)).

  When considering a flat-curved cavity with cavity length L, dx and dt can be expressed as;

(dx)    1  ( L*g     L ) (a2)
(  ) = --- (           )*(  )
(dt)   1-g ( -(1-g)  0 ) (a1)


using misalignments of mirrors(a1,a2). Here, mirror1 is curved, and mirror2 is flat. See Kakeru document /users/OLD/kakeru/oplev_calibration/oplev.pdf for derivation.

  So, power decrease by flat mirror misalignment can be expressed as

dP/P = pi*L/lamb * g/(1-g)/sqrt(g*(1-g)) * a2**2

  For curved mirror is

dP/P = pi*L/lamb * 1/(1-g)/sqrt(g*(1-g)) * a1**2

  We can derive g-factor by measuring dP dependance on a1/a2.


Script:
  My script lives in /opt/rtcds/caltech/c1/scripts/dither/gfactormeasurement/plotgfactor.py.
  It least fitts data with parabola (scipy.optimize.leastsq) and gets g-factor value from bisection (scipy.optimize.bisect).


Result:
  Below are the plots of fitted curves.

ITMY_YAW_20130305_DC.pngPRM_PIT_20130305a_DC.pngPRM_YAW_20130305b_DC.png


Systematic effect:
  [oplev calibration] We noticed QPD rotation when calibration oplevs (elog #8232). ~5 deg of rotation makes 10% of systematic error to the oplev calibration and this introduces ~0.04 of error to g-factor values. This

  [oplev linear range] Oplev linear range is ~100 urad, so this is OK.

  [assumption of flat PR2/PR3] Result here doesn't tell you g-factor of PRM itself, but some "effective g-factor" of PRM/PR2/PR3 combination. We can compare with FINESSE result.

  [intra-cavity power drift] If there's significant intra-cavity power drift during the measurement, if effects parabola fitting. We can make this affect small by sweeping the mirror alignment in both direction and take average.


By the way:
  I kept getting PRC g-factor of something like 0.999999 because I had power normalization mistake in my calculation. My script worked for Yarm because TRY is already normalized.
  Also, I was multiplying the oplev calibration factor wrong last night (see elog #8230).

Next:
  - Compare with FINESSE result.
  - Is this g-factor enough? Is this presicion enough? Calculate from mirror angluar motion.
  - More stable lock of PRMI.
  - Try dithering method to measure g-factor to check consistency and also to study systematic effect.

  8234   Tue Mar 5 18:36:27 2013 JenneUpdateLSCLSC whitening triggering started

More effort at debugging the LSC whitening. 

Today I tried moving things over to the c1tst model, which runs on the y-end computer, but I crashed c1iscey.  I rebuilt the TST model to a known good state, then cycled the power on c1iscey, and the computer came back up fine. 

I have now backed off and am just writing the code inside a little wrapper script, so that I can just compile and test the code completely independent from the realtime system.  Then once I get all the bugs out, I can try again installing on the actual system.

Still, there are no changes to the functionality of the c1lsc model.  There will not be until I get the c-code for matrix parsing debugged.

The logic, in non-diagram form (I'll make a diagram, but so you can read without waiting):

*** C-code

* Inputs is an array of degree of freedom triggers, the same schmidt triggers used for main LSC locking. (This means it also uses the same thresholds as main triggers.  Side note, now that the WAIT command (see below) works, I want to change the filter module triggers to use the same main trigger, and then just wait a specified time before turning on.)

* Parse the LSC input matrix (internal to the c-code).

     * This tells you which photodiode is being used to control which degree of freedom.

* Multiply rows of the LSC input matrix by the degree of freedom triggers (the same trigger as the main LSC triggers, which is a schmidt trigger).

     * This gives a matrix, where non-zero elements indicate that a photodiode is supposed to be used for a degree of freedom, AND that DoF has been triggered (is locked or has flashed).

* Sum along the columns of the matrix.

     * If a column has a non-zero element, that means that that PD quadrature is used, and has been triggered (by any DoF).

* Apply OR to I and Q quadratures of each PD. 

      * Since the phase rotation happens after whitening and dewhitening, if either I_ERR or Q_ERR is requested (used and triggered), we need to turn on the whitening for both channels.  I am hopeful that this doesn't cause problems for cases when we want to use both quadratures of a PD to control 2 degrees of freedom, but I haven't yet put much thought into it.  COMMENTS WELCOME on this point.

*  Output of c-block is array of PD triggers.  So if either AS11I or AS11Q is triggered, output a "1" for the first element, which corresponds to AS11, etc.

*** LSC model

* Give GoTo/From flag for each DoF trigger to the mux of inputs.

* Go through c-code

* Demux outputs into GoTo/From flags, one per PD (one flag for AS11, one for AS55, and one for ASDC...DC elements count separately, even though they're derived from the same physical PD).

* For each PD, trigger flag goes through WAIT c-code

   * This allows you to define a wait time, in seconds, with an EPICS variable. 

   * Starts counting the wait time as soon as it receives a "1".  Resets counter each time it receives a "0".

    * Output of wait function is ANDed with the current (non-delayed) trigger.

         * This allows for cavity to flash, but if it's not still locked after the wait time, don't actually flip any switches.

* Use delayed ANDed trigger to flip the FM1 switch on both the I and Q filterbank for that PD.

  8233   Tue Mar 5 17:23:06 2013 ChloeUpdate QPD Adding/Subtracting Circuit

Today I finished building the adding/subtracting circuit for the QPD and tested that the QPD could see a laser moving across its visual field for both pitch and yaw. It didn't seem to behave weirdly (saturate) at the edges, but I need to test this more carefully to be sure.

However, this circuit uses many op amps, which will cause problems for building the actual circuit to fit into the QPD box. I am trying to figure out how to do this with fewer op amps (both with a quad op amp for amplifying the signals from the QPD and by summing/subtracting the signals with a single op amp instead of 3).

I finally got around to asking Steve to order more breadboards! Trying to determine what would be a good QPD to order for the final circuit, since we do not have any unmounted QPDs that aren't ancient. I'll read up on things I don't know enough about (namely op amps).

  8232   Tue Mar 5 17:09:30 2013 yutaUpdateSUSoplev calibration for ITMY

[Manasa, Sendhil, Yuta]

We calibrated oplev for ITMY. Calibration factor for C1:SUS-ITMY_OL(PIT|YAW)_IN1 are;
  OLPIT: 6.29 +/- 0.11 counts/mrad
 
OLYAW: 5.74 +/- 0.09 counts/mrad

Note that there was ~10% of coupling between pitch and yaw.
This is large considering statistical error, but I think this is from QPD mounted rotated (by ~5 deg).

Method:
  Same as in elog #8221.

Result:
 
moved in y: ITMY_PIT.png      moved in x: ITMY_YAW.png

micrometer    OLPIT          OLYAW
moved in y    3.14 +/- 0.05  0.27 +/- 0.03
moved in x   -2.87 +/- 0.04  0.17 +/- 0.03    (counts/mm)


  Measured the path length of ITMY oplev returning beam was 2000 +/- 20 mm. This gives you the calibration factors above.

  ~10 % coupling between OLPIT and OLYAW can be explained by QPD rotation by ~ 5 deg, which seems not unreasonable.

  8231   Tue Mar 5 14:59:32 2013 SteveHowToVACdry-forepump replaced

TP3 turbo pump's dry-foreline pump was replaced.

How to do it:

The pump should be replaced when it's performance  <1.0 - 1.3 Torr

Set up valve configuration as shown at Atm1: close in this order VAEE, VASV, VABS, VASV, VASE, VA6 and V5,

Turn TP3 off at it's controller in the rack. Wait till is stops, so you can read 760 Torr at TP3 foreline gauge

Disconnect intake, exhaust vacuum seals and replace pump. Reconnect vacuum fitting and start it up.

Confirm operational details on the front of the controller: 50 K_RPM, 0.2A and <100 mTorr

Reset valves in reverse order

PS: the average life of the tip seal on the Varian SSH-110 dry-pump is about 1 year

      This pump " ser LP1007L556 " seal  made new record of 668 days: thanks to Bob Taylor who is replacing these seals

 

Attachment 1: changingforepump.png
changingforepump.png
Attachment 2: TP3wFp600.jpg
TP3wFp600.jpg
  8230   Tue Mar 5 06:27:14 2013 yutaUpdateLSCintra-cavity power dependence on mirror misalignment

I measured intra-cavity power dependance on mirror misalignment.
Intra-cavity power of PRC in PRMI degrates roughly 20 % when there's 0.5 mrad 5 urad misalignment. (edited by YM)
Currently, PRMI lock is not so stable, so it is hard to do this measurement and error bars are huge.

Measurement method:
  0. Align the cavity and lock.
  1. Misalign one optic and measure oplev output value and intra-cavity power.
  2. Also, dither the optic in pitch or yaw in 8.5 Hz and get demodulated amplitudes at 8.5 Hz of oplev output and intra-cavity power using tdsdmd.
  3. Misalign the optic again and do the same things.

  1. gives intra-cavity power dependence on mirror misalignment directly, but 2. should give better S/N because of dithering.


Scripts:
  /opt/rtcds/caltech/c1/scripts/dither/dithergfactor.py does these things, and ./plotgfactor.py plots the result.
  They work quite well, but it should be made better so that

  - it checks if the cavity is locked
  - automatically change the oplev calibration factor for each optic
  - automatically adjusts the region and modulation amplitude
  - read data with better error evaluation

  etc...


PRMI alignment:
  Y green looks like it drifted quite a lot somehow. If we start aligning Yarm to Y green, we get AS and POP beam at different spot on camera compared with last week. Also, TRY and TRX only goes as high as ~0.7. Since we have A2L now (elog #8229), let's start using Yarm spot positions as input pointing reference.


PRMI locking details:
  Same as in elog #8212, but I changed gains in the lock acquisition mode.

  == PRMI carrier ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg,  MICH_GAIN = -0.2, feedback to ITMX(-1),ITMY(+1)
  PRCL: REFL55_I_ERR, REFL55_PHASE_R = 70 deg, PRCL_GAIN = 1.0, feedback to PRM

  I made gainx5 in LSC_MICH filter bank so that it increases the overall gain when locked by trigger.
  I also made gainx0.3 in LSC_PRCL filter bank so that it reduces the overall gain when locked by trigger.


Result for PRC in PRMI:
  For PRMI, I couldn't done dithering method because dithering takes time to measure and I could not hold PRMI locking during the measurement.
  Below is the result when reading just the DC values. Mirror angle is calibrated by oplev (elog #8221). Error bars are huge because of beam motion mainly in yaw.


PRM in pitch: PRM_PIT_20130305a.png    PRM in yaw:PRM_YAW_20130305b.png


Results for the arms:
  For the arms, I could do both in DC and dithering. Below are the results, but ITMs misalignments are not calibrated because we don't have calibrated oplev yet.
  Results for the arms can be used to verify this method because we know g-factors of the arms from mode scan.


ITMX in yaw: ITMX_YAW_20130305.png    ITMY in yaw: ITMY_YAW_20130305.png



By the way:
  I found C1:SUS-ITMY_LSC_GAIN is somehow set to be 2.895 recently. I think this should be 1.0. Maybe this is why we had actuation imbalance in ITMs(elog #8212).


Next:
 - more stable lock
 - calibrate ITM oplevs to apply this method to the arms
 - derive g-factor from these measurements
 - measure PRM angular motion spectra using calibrated oplev

  8229   Tue Mar 5 01:43:04 2013 JenneUpdateASSArm A2L measurement script finished

In either .../scripts/XARM or ...../scripts/YARM run either A2L_XARM or A2L_YARM.

The wrapper script will, like the MC script, open a striptool so you can monitor the lockin outputs, setup the measurement, run the measurement, including misbalancing coils on the optics for calibration, and then calculates the spot positions.  It records the measurement in a log file in /data_spotMeasurements under each arm's directory.  The wrapper script then runs the plotting script which reads the logfile, and plots all past measurements.

Here is that plot for the Yarm:

YARMdecenter.png

The first two points were measured within a few minutes of eachother, the third set of points was after input pointing adjustment during IFO alignment.  Clearly the pointing that optimized the cavity transmission (trying to leave the test mass mirrors alone, and only moving TT1 and TT2) does not also give the best spot centering.  I claim that this is a result of the arm being aligned to the green beam, which was never locked to the 00 mode when we were at air.  This is a lesson learned....take the time to deal with the green beams.

  8228   Tue Mar 5 00:24:52 2013 ManasaConfiguration40m UpgradingEndtable upgrade for auxiliary green laser : customized mount

 Drawing of customized mount for endtable doubler crystal (PV40+PVP2+NP9071) modified and updated on the endtable upgrade wiki page.

  8227   Mon Mar 4 21:05:49 2013 Max HortonUpdateSummary PagesMultiprocessing and Crontab

Multiprocessing:  In its current form, the code uses multiprocessing to the maximal extent possible.  It takes roughly 2600 seconds to run (times may vary depending on what else megatron is running, etc.).  Multiprocessing is only used on the process_data() function calls, because this by far takes the longest.  The other function calls after the process_data() calls take a combined ~120 seconds.  See http://nodus.ligo.caltech.edu:8080/40m/8218 for details on the use of Multiprocessing to call process_data().

Crontab:  I also updated the crontab in an attempt to fix the problem where data is only displayed until 5PM.  Recall that previously (http://nodus.ligo.caltech.edu:8080/40m/8098) I found that the crontab wasn't even calling the summary_pages.py script after 5PM.  I changed it then to be called at 11:59PM, which also didn't work because of the day change after midnight.

I decided it would be easiest to just call the function on the previous day's data at 12:01AM the next day.  So, I changed the crontab.

Previous Crontab:

59 5,11,17,23 * * * /users/public_html/40m-summary/bin/c1_summary_page.sh 2>&1

New Crontab:

0 6,12,18 * * * /users/public_html/40m-summary/bin/c1_summary_page.sh 2>&1
1 0 * * * /users/public_html/40m-summary/bin/c1_summary_page.sh $(date "+%Y/%m/%d" --date="1 days ago") 2>&1

For some reason, as of 9:00PM today (March 4, 2013) I still don't see any data up, even though the change to the crontab was made on February 28.  Even more bizarre is the fact that data is present for March 1-3.  Perhaps some error was introduced into the code somehow, or I don't understand how crontab does its job.  I will look into this now.

Next:

Once I fix the above problem, I will begin refactoring the code into different well-documented classes.

  8226   Mon Mar 4 20:03:42 2013 AnnalisaHowToCOMSOL TipsStudy of mirror mount eigenfrequencies

 I studied the eigenfrequencies of a mirror mount designed with COMSOL.

I imposed fixed constraints for the base screws and for the screw connecting the base with the pedestal. Note that the central screw is connected to the base only for a small thickness, and the pedestal touches the base only with a thin annulus. This is in way to make a better model of the actual stress.

Shown in fig. 2 is the lowest eigenfrequency of the mount.

I' going to change the base and study the way the eigenfrequency vary, in way to find the configuration which minimizes the lowest eigenfrequency.

 

Attachment 1: MirrorSupport1.png
MirrorSupport1.png
Attachment 2: MirrorSupportEig1.png
MirrorSupportEig1.png
Attachment 3: pedestal.png
pedestal.png
Attachment 4: Base2.png
Base2.png
  8225   Mon Mar 4 19:52:03 2013 Jamie, YutaUpdateGeneralinput pointing mirror (TT1/TT2) control improved

We improved the active tip-tilt (TT) controllers such that they now have filter banks at the PIT/YAW inputs, and at the coil outputs:

TT.png

This allowed us to do a couple of things:

  • normalize the matrix to unity, and move overall gains into the filters (we moved x100 gain into PIT/YAW)
  • slider control is now PIT/YAW OFFSET
  • potentially do coil balancing
  • allow for input excitiations
  • automatically record EPICS values

These are all big improvements.  The TT MEDM screens were appropriately updated.

We had to rebuild/restart c1ass, which reset the TT pointing.  We recorded all the values before hand and were able to recover the pointing easily.  Interestingly, there did appear to be hysteresis in pitch, which is maybe not entirely unexpected, but still worth nothing.

 

  8224   Mon Mar 4 19:38:21 2013 CharlesUpdateGeneralIntensity Stabilization and Control Systems

 I have been studying Jamies master's thesis concerning intensity stabilization of a solid-state laser (the 1064 nm specifically) to the ~10^-9 level, as well as relevant supporting material. I have also been reading about general control systems, photodiodes and acousto-optic modulators to help facilitate work on the ISS. 

Now that Altium has been properly installed, I have also begun familiarizing myself with the program and general libraries of boards and devices that have already been modeled with the program.

  8223   Mon Mar 4 18:11:10 2013 JamieUpdateSUSCleaning up suspension POS inputs

I did a little bit of cleanup of the suspension POS inputs today, both in model and MEDM land.

model

sus_single_control.mdl was simplified such that now there are just two position inputs:

  • POS: with LSC filter
  • ALTPOS: unfiltered

The regular LSC inputs go into POS, and any optic-specific extra pos inputs go into ALTPOS after being properly filtered and summed.

So for instance, MC2_MCL and MC2_ALS are filtered and summed then go into MC2 ALTPOS.  The ETM ALS inputs go into ALTPOS.

I modified the GLOBAL DAMPING outputs so that they are filtered in the GLOBAL block before being sent to be summed before going into ALTPOS for {I,E}TM{X,Y}.

All suspension models were rebuilt/installed/restarted.

MEDM

The SUS_SINGLE.adl template screen was modified such that the POS button now points to optic-specific POS filter screens at:

/opt/rtcds/caltech/c1/medm/master/sus/SUS_$(OPTIC)_POS.adl

For MC1, MC3, PRM, BS, SRM these are links to SUS_SINGLE_POS.adl.  The rest of the suspensions (MC2, {I,E}TM{X,Y}) now have custom screens that are variations of SUS_SINGLE_POS but with their extra filter screens added in.  For instance, here is the new SUS_ETMX_POS.adl:

SUS_ETMX_POS.png

This gets rid of all the white screen crap that was in here before.

All of this has been committed to the SVN.  NOTE: symlinks were heavily used when sorting this stuff out, so check for symlinks when modifying in the future.

  8222   Mon Mar 4 17:32:26 2013 yutaBureaucracyGeneralflowchart for PRMI g-factor measurement

I made a very useful flowchart for the week. Our goal for the week is to measure g-factor of PRC in PRMI.

PRMIgfactorPlan.png

  8221   Mon Mar 4 16:46:31 2013 yutaUpdateSUSoplev calibration for PRM

[Manasa, Sendhil, Yuta]

We calibrated oplev for PRM. Calibration factor for C1:SUS-PRM_OL(PIT|YAW)_IN1 are;
  OLPIT: 15.6 +/- 0.3 counts/mrad
  OLYAW: 17.8 +/- 0.3 counts/mrad


We needed these values for g-factor measurement of PRC using angle dithering method.

What we did:
  1. Adjusted QPD offsets (C1:SUS-PRM_OL[1-4]_OFFSET) by zeroing the output when turned oplev laser was turned off.
  2. Centered PRM oplev beam on QPD using steering mirror.
  3. Mounted PRM oplev QPD on a xy-stage and centered the beam on QPD.
  4. Moved QPD in x and y using micrometers and measured dependance of C1:SUS-PRM_OL(PIT|YAW)_IN1 on QPD displacement.
  5. Measured the path length of PRM oplev returning beam. We get the in-vac path length using optical layout CAD. We measured out of vac path using scale and tape measure.
  6. Dismounted PRM QPD from the stage and put it back to the original position.

Result:
  1. Figures below are OLPIT/OLYAW dependance on micrometer displacement moved in x and y. Error bars are measured fluctuation in the signal.


moved in x:PRM_PIT.png       moved in y:PRM_YAW.png


  2. We fitted the result by error function to get slope at zero crossing point. We also linear-fitted the other degree of freedom to get cross coupling between x and y. Slopes we get were;

micrometer    OLPIT          OLYAW
moved in x    4.68 +/- 0.08  0.01 +/- 0.03
moved in y   -5.32 +/- 0.10  0.11 +/- 0.03    (counts/mm)


  3. Measured the path length of PRM oplev returning beam was 3340 +/- 20 mm. This gives you the calibration factors above.

Discussion:
  [Precision] According to Jamie's calculation, expected g-factor for PRC in PR2-flipped PRMI is 0.966/0.939 (elog #8068). We care about the g-factor change in ~0.01. Also, intra-cavity power dependance on mirror angle is proportional to 1/(1-g). So, we need to calibrate mirror angle in ~few 10 % of precision in order to get useful g-factor from angle dithering measurement. Measurement precision here meets this requirement.

  [x/y coupling] Measured x/y coupling was less than 2 %. We assumed gains in 4 QPD quadrants are same, and assumed QPD is mounted well in x/y axes. These assumptions are OK considering precision we need.

  [x/y difference] Calibration factors for OLPIT/OLYAW are different by ~10 %. This is not so crazy considering the incident angle on the QPD (~20 deg) and elliptic beam shape.

  8220   Mon Mar 4 16:26:45 2013 BrettUpdateSUSGlobal Damping Noise Measurement

Here is an amplitude spectrum plot of y-arm cavity noise with a 50 Hz cutoff damping filter of the form zpk(0,[50;50],1). The low passing of this filter was intentionally extremely poor in order to see the damping noise in the cavity. The blue trace is the noise with no damping, which may be considered the 'best case' scenario from a noise point of view. The green has regular local damping on the ITMY. The ETMY has no damping for this measurement because the cavity control feedback to the ETMY takes care of its control when the cavity is locked. Notice the the large increase in noise from 40 Hz to 250 Hz, up to 1 order of magnitude. This noise is from the OSEM sensors passing through the damping loops. The red curve shows the y-arm noise with the exact same damping, except it is now applied in the global scheme. In this case, the damping noise falls completely below the baseline level of the cavity and becomes indistinguishable from the 'no damping' case.

If the damping injected enough noise I'd expect we would see a drop of 50 to 80 times switching from local to global. That is, the same factor measured in the transfer functions listed in log entry 8193.  However, the damping noise is only at most 1 order of magnitude above the baseline in this measurement. We would have to increase the damping noise by about another order of magnitude before we could expect to see the global damping noise in the cavity measurement.

The units of the cavity displacement in the plot were calculated using the 1.4e12 counts per meter calibration in log 6834. The measured UGF of the LSC loop at the time was 205 Hz. The peak in the plot above 200 Hz appears to be from this unity crossing. Moving the UGF also moves this peak.

Moral of the story: global damping can isolate the damping noise pretty well from the cavity signal.

Attachment 1: YARM_Noise.png
YARM_Noise.png
  8219   Mon Mar 4 11:30:47 2013 SteveUpdatePEMair cond problem

 

 The air cond is out of order at the area covered by racks 1 X 1  through 1 X 7

The arm X and Y AC units are working.

Attachment 1: 40dAC.png
40dAC.png
  8218   Mon Mar 4 10:41:18 2013 Max HortonUpdateSummary PagesMultiprocessing Implementation

Update:

Upon investigation, the other methods besides process_data() take almost no time at all to run, by comparison.  The process_data() method takes roughly 2521 seconds to run using Multiprocessing with eight threads.  After its execution, the rest of the program only takes 120 seconds to run.  So, since I still need to restructure the code, I won't bother adding multiprocessing to these other methods yet, since it won't significantly improve speed (and any improvements might not be compatible with how I restructure the code).  For now, the code is just about as efficient as it can be (in terms of multiprocessing).  Further improvements may or may not be gained when the code is restructured.

  8217   Mon Mar 4 09:55:33 2013 ranaUpdateLSCLSC whitening triggering started

  How about posting a logic flow diagram? Is the idea to trigger only on the power signals to determine the lock state? Is the hysteresis going to be done in the same way as the main filter bank triggers?

  8216   Mon Mar 4 09:51:26 2013 SteveUpdateASSIP-ANG is still not real

 IP-ANG  is coming out of ETMY without clipping. The beam is very high on the pick off mirror at the end table but it is still missing the qpd .

 

Attachment 1: IP-ANG.png
IP-ANG.png
  8215   Sun Mar 3 22:16:46 2013 JenneUpdateLSCLSC whitening triggering started

[Jenne, Annalisa]

We have started working on writing the c-code to parse the LSC input matrix, to see which PD is used for what degree of freedom, and to output a trigger for the PD.  The code is in ..../isc/c1/src, and there is a little block in the LSC code to the left of the triggering stuff.  Right now, the output of the c-code just goes to some temporary EPICS channels, so that we can see if things are working before we actually implement it.  At this time, there is no change to how the LSC model runs.

I can't figure out a bug in my c-code though.  Right now it's all commented out, so that the LSC model would start, but if I try to sum all of the elements in an array, the model compiles fine, but it won't start running.  I'm going to ask Jamie about it tomorrow.  I have a less-tidy backup plan if we can't get this figured out.

If I have time on the IFO to check that this works tomorrow, I expect another few hours of work (2?  3?), and then we'll have whitening filter triggering.

  8214   Sat Mar 2 20:09:12 2013 yutaUpdateRF Systemphase tracker: DAQ noise limited

[Koji, Yuta]

We found that our phase tracker noise is currently limited by the noise introduced in DAQ.
We confirmed that the frequency noise was improved from 2 Hz/rtHz to 0.4 Hz/rtHz by increasing the gain of the whitening filter.
The whitening filters should definitely be refined.

What we did:
  1. Put constant frequency RF input to the beatbox from Marconi and measured noise spectrum of the beatbox output(BEATX I) after the whitening filter with a spectrum analyzer. Noise floor level was ~0.2 Hz/rtHz at carrier frequency range of 15-100 MHz. Calibration factor of the beatbox output was ~380 mV/MHz.

  2. Measured noise spectrum of C1:ALS-BEATX_FINE_I_OUTPUT(figure below). The noise floor didn't change when there was RF input of 100 MHz from Marconi(blue) and DAQ input was terminated (green). Also, C1:ALS-BEATX_FINE_I_IN1(which is before unwhitening filter) showed a flat spectrum. These show our spectrum is limited by DAQ noise, which is introduced after the whitening filter.

  3. We increased the gain of whitening filter by x20 to show frequency noise performance can be improved by better whitening filter(red). But we can not use this setup as the other quadrature will be saturated by a too much gain at DC. Thus we need to carefully consider the signal level and the gain distribution of the whitening filters.
frequencynoisewhitening.png


Next:
  - Better whitening filters. The current one consists of zero 1 Hz and pole 10 Hz with DC gain of 5 using SR560.
  - Better beatbox. We can increase the RF input power to the mixer and unify the preamplifier and the whitening filter in the box.

  8213   Sat Mar 2 14:52:02 2013 ranaUpdateLSCstable lock of PRMI

  Whereas the sensing matrix coefficients for ITMX/ITMY in REFL_I indicate an actuation imbalance, the disparity in the ITMX/ITMY to AS_Q elements does not. However, they do indicate why there is a PRM to AS_Q coupling at all.

I would recommend setting up the triggering so that the REFL & AS whitening is turned on AFTER lock acquisition and using a frequency of ~100-300 Hz for the sensing matrix measurement to fix these issues.

  8212   Sat Mar 2 05:53:15 2013 yutaUpdateLSCstable lock of PRMI

I tuned alignment, gains and filters to achieve stable lock of PRMI.
It now locks quite stably with UGF of ~100 Hz. Measured power recycling gain at maximum is ~ 25.

Locking details:
  == PRMI carrier ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg,  MICH_GAIN = -1, feedback to ITMX(-1),ITMY(+1)
  PRCL: REFL55_I_ERR, REFL55_PHASE_R = 70 deg, PRCL_GAIN = 0.3, feedback to PRM

  MICH servo is always on. PRCL loop turns on by trigger using POP DC. Boost filters and resonant gains turn on by triggers using POP DC.
  Gain normalization was not used.


Openloop transferfunctions:

  MICH: UGF ~90 Hz, phase margin ~40 deg
  PRCL: UGF ~100 Hz, phase margin ~50 deg (cf. Fitted gain was same as half-PRC: elog #8053)
LSCMICHOLTF_PRMI.png    LSCPRCLOLTF_PRMI.png



Power recycling gain:

  POP DC when unlocked is 6, when locked is 2200-2500, and when dark is 0. So, power recycling gain is ~ 22 to 25. Value without any loss in PRMI is 45 (elog #6947). Alignment was pretty critical to achieve this recycling gain and stable lock.
  There was oscillation at 630 Hz when locked, which is similar to the one we saw in POX11 (elog #8203).


Youtube:





AS(top left), POP(top right), REFL(bottom left), and ETMYT(bottom right). ETMY was mis-aligned, but you can see the beam at ETMYT after PRMI was carrier locked.



MICH/PRCL coupling:

  I measured "sensing matrix" of PRMI by tickling PRM/ITMs/BS at 8.5 Hz and measuring 8.5 Hz peak height of AS55_Q, REFL55_I spectra during PRMI lock (attached is an example measurement of PRM). Below table is the result. AS55_Q has ~5% of sensitivity to PRCL compared with MICH. Also, BS introduces REFL55_I signal considerably. And also, there seems to be an imbalance in actuation efficiency between ITMX and ITMY.

actuation AS55_Q_ERR   REFL55_I_ERR
ITMX      +11.4        +0.80
ITMY      +33.0        +1.06
BS        +50.8        +1.90
PRM       - 0.7        +1.05



AS clipping:
  AS was clipped inside the vaccuum the other day(elog #8198). So, I tried to avoid AS clipping by aligning BS this morning. But it turned out that avoiding AS clipping by BS makes ITMX beam centering worse. Maybe we should center the beam on Yarm first next week.


Next:
 - calculate expected PRMI recycling gain with loss, PR2 filpped
 - beam centering on the arms
    - IPANG, IPPOS, Y green, X green
    - PRMI g-factor measurement

Attachment 3: PRMtoAS55REFL55.png
PRMtoAS55REFL55.png
  8211   Sat Mar 2 00:23:19 2013 ranaSummaryCOCPhase Maps measured of the ATF flat mirrors

I took the two 'flat' 2" mirrors over to Downs and Garilynn showed me how to measure them with the old Wyko machine.

The files are now loaded onto our Dropbox folder - analysis in process. From eyeball, it seems as if the RoCs are in the neighborhood of ~5 km, with the local perturbations giving ~10-15 km of curvature depending upon position (few nm of sage over ~1 cm scales)

Koji's matlab code should be able to give somewhat more quantitative answers...

Ed: Here you are. "0966" looks good. It has RoC of ~4km. "0997" has a big structure at the middle. The bump is 10nmPV (KA)

 

Attachment 1: 0966_0997_phasemap.pdf
0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf 0966_0997_phasemap.pdf
  8210   Sat Mar 2 00:09:31 2013 ranaUpdateLSCXarm oscillation stopped

  Don't use resonant gain - it can lead to a loop instability since it makes the loop have 3 UGFs.

Just use a elliptic bandstop filter at this harmonic frequency separately for each test mass. There are many detailed examples of this in elog entries from Rob and I over the past ~10 years. This bandstop should get clicked on automatically after lock acquisition.

  8209   Fri Mar 1 18:23:28 2013 JamieUpdateComputer Scripts / Programsupdated version of "getdata"

I updated the getdata script so that it can now handle downloading long stretches of data.

  /opt/rtcds/caltech/c1/scripts/general/getdata

It now writes the data to disk incrementally while it's downloading from the server, so it doesn't fill up memory.

I also added a couple new options:

* --append allows for appending to existing data files

* --noplot suppresses plotting during download

  8208   Fri Mar 1 16:58:37 2013 yutaUpdateLSCXarm oscillation stopped

POX11 oscillation at 630 Hz was stopped by installing 630 Hz resonant gain to LSC_XARM.
After few hours, oscillation stopped. So I removed the resonant gain.
Our guess is that 630 Hz peak is some violin mode or something, and it was excited somehow, and didn't stopped for very long time because of its high Q. It coupled into POY11 somehow (scattering, electronics, etc).

Attachment 1: POX11_630Hz.png
POX11_630Hz.png
  8207   Fri Mar 1 16:37:45 2013 BrettUpdateSUSGlobal Damping Update

Brett and Kamal

The global damping testing for the week is now complete. The c1sus.mdl simulink diagram settled on the attached screenshot. The top level of c1sus.mdl is shown on the left zoomed in over the new global damping block. The right shows the inside of that block. Also attached in the second screenshot are two of the modal damping MEDM screens. The left shows the main overview screen, the right shows the global damping filters. The overview screen is called C1SUS_GLOBAL.adl and is found in ...medm/c1sus/master/.

We have measured transfer functions and power spectra that show that global damping, with just a moderate amount of tuning (30 minutes of work) reduces the OSEM damping noise seen by YARM_IN1 by a factor between 50 and 80. Log 8193 highlights the transfer function measurements. The power spectra directly measure the noise in the cavity. I am not putting that data here because I have to catch. I will process the data and post it here later.

Overall the global damping tests appear to have been successful, isolating (not removing) the test mass damping noise from the cavity by almost 2 orders of magnitude. Presumably even more isolation is possible with more tuning.

Attachment 1: GlobalDamp_Simulink.png
GlobalDamp_Simulink.png
Attachment 2: GlobalDampScreens.png
GlobalDampScreens.png
  8206   Fri Mar 1 14:50:36 2013 SteveUpdateVACpumpdown day 8

 Atm1, RGA background scan with TP3 pumping at day 7

Atm2,  IFO  CC1 = 1.1e-5 Torr at day 8

 Vacuum valve VM3 closed and VM2 opened, it switched the RGA to IFO scanning mode. CC4 = 3.5e-6 Torr

TP3 is pumping on the annuloses, their pressures are 1-3 mTorr

Vacuum condition reached the state of  Vacuum Normal. The MEDM screen "Current State" channel does not work because we disconnected some valves after the VME swap CAOS

Note: TP1 Maglev performance on Atm2  CC4 = 4e-9 Torr when VM2 valve was open

 

Attachment 1: bgTP3d7.png
bgTP3d7.png
Attachment 2: RGAonIFO.png
RGAonIFO.png
  8205   Fri Mar 1 08:19:40 2013 SteveUpdateASSdrift

 Early morning drift in pitch. This plot is meaningless because there is no real  light on IP-Ang

The beam is clipping on the pick off mirror at ETMY chamber. The beam is half beam size too high. Yaw is perfect

 

Attachment 1: drift.png
drift.png
  8204   Fri Mar 1 02:49:34 2013 JenneUpdateASSArm A2L measurement

I haven't finished debugging the scripts so that the measurement is fully automatic, like the MC, but I did measure the arm spot positions just now. 

These numbers aren't especially precise....I just picked numbers off of a StripTool plot, but they give us a good idea of how very far off we are.  Also, I don't know yet which way the signs go...I have to think about that in terms of the direction I mis-balanced the coils.  It's the same convention as the MC though.  You can see in the attached quad camera image (quadrants match the corners of the table) that these numbers aren't unreasonable.

ETMY   ETMX  
Pit 4 mm Pit 4 mm
Yaw -1.5 mm Yaw 6 mm
ITMY   ITMX  
Pit -3 mm Pit -3 mm
Yaw 4 mm Yaw -4 mm

QUAD1_1046168242.bmp

EDIT: It occurs to me now, a little later, that it had been at least half an hour since I last realigned the cavities, so some of this apparent miscentering is due to the input pointing drift. That doesn't account for all of it though. Even when the cavities have very high transmitted power, the spots are visibly miscentered.

  8203   Fri Mar 1 01:17:06 2013 JenneUpdateLSCXarm oscillation

There is an oscillation in the Xarm at 631Hz, which is not in the Yarm.  There is a small peak in POY11_I at this frequency, but only when the Xarm is locked.  If the Xarm unlocks, the peak disappears from POY.  The peak is 3 orders of magnitude larger in POX than in POY, and 4 orders of magnitude larger than the POY noise when this peak is not present.  In the plot, I have turned off the POY whitening, so that its situation is the same as POX (we still need to fix POX whitening switching).  Dark noise (MC unlocked) is the same for both PDs.

POX11_630Hz_osc.pdf

  8202   Thu Feb 28 14:51:43 2013 JenneUpdatePEM"Rock Monster" causing low frequency noise

Manasa and I have been wondering why the low frequency seismic noise has been different in the last few days/weeks.  I went over to visit the Rock Monster, a.k.a. the Earth Surface Dynamics Laboratory next door, and the grad student turned it up for me (increased water flow significantly) for a few minutes.  Then I came back to look at the BLRMS, and you can see the low frequency increase a few minutes ago, when he turned it up on high.  He said that they run it on low almost all the time, 9am-5pm ish, and on high for an hour or so at a time periodically during the day.

RockMonsterOn_ForFewMinutes.png

  8201   Thu Feb 28 14:19:20 2013 Max HortonUpdateSummary PagesMultiprocessing Implementation

Okay, more memory would definitely be good.  I don't think I have been using MEDM (which Jamie tells me is the controls interface) so making a cronjob would probably be a good idea.

  8200   Thu Feb 28 06:51:17 2013 yutaUpdateRF Systemphase tracker: noise level

I measured noise level of the phase tracker by inputting constant frequency RF signal from marconi.
Measured frequency noise was ~2 Hz/rtHz @ 100 Hz. It's not so good.

What I did:
  1. Unplugged 11MHz marconi and put RF signal to the beatbox from this. Frequency and amplitude I put are 100 MHz and -3 dBm.
  2. Measured spectra of phase tracker outputs, C1:ALS-BEATX_FINE_PHASE_OUT, C1:ALS-BEATY_FINE_PHASE_OUT.
  3. Calibrated using the factor I measured (elog #8199).
  4. Put marconi back to orignal settings.

Result:
frequencynoise.png

Discussion:
  - According to Schilt et al., this noise level is not so good.
  - By changing the delay-line cable length or optimizing whitening filter etc., we can improve this.
 

  8199   Thu Feb 28 05:54:54 2013 yutaUpdateRF Systemphase tracker: calibration

I swept the frequency of RF input to the beatbox to calibrate and check linearity range of phase tracker.
Calibration factors are;
  C1:ALS-BEATX_FINE_PHASE_OUT    52.1643 +/- 0.0003 deg/MHz
  C1:ALS-BEATY_FINE_PHASE_OUT    51.4788 +/- 0.0003 deg/MHz


There was systematic error to the linearity check, but at least, calibration factor changes less than 50 % in the frequency range of 10 MHz to more than 500 MHz.


What I did:
  Used network analyzer(Aligent 4395A) to sweep the frequency RF input to the beatbox and getdata of phase tracker signal. I swept from 10 Hz to 500 MHz with 501 points in 50 sec. This sweep is slow enough considering we could lock the 40m arms (typical speed of a mirror is 1 um/s, so bandwidth of the phase tracker should be more than 1 um/sec / 40 m * 3e14 Hz = 75 MHz/s).
  RF amplitude was set to be -3 dBm and splitted into BEATX and BEATY.


Result:
  Plots for BEATX and BEATY are below;
ALS-BEATX_FINE_PHASE_OUT.pngALS-BEATY_FINE_PHASE_OUT.png


Discussion:
  - Considering delay line length is ~30m, expected calibration factor is;

    2*pi*l/v = 2*pi * 30 m / (2e8 m/s) = 0.94 rad/MHz = 54 deg/MHz

so, this calibration is reasonable.

  - Since frequency sweep of network analyzer is not continuous, phase tracker output is like steps with some ringdown. This makes some systematic error for checking linearity. I'm planning to do slower sweep or continuous sweep. Also, the phase tracker seems like he can exceed 500 MHz.

  8198   Thu Feb 28 03:41:31 2013 KojiUpdateLSCPR gain ~ 25 from PRMI carrier lock

VERY GOOD!
This is how the carrier lock PRMI should look like.

- There is more room to improve the differential ITM alignment to make the dark port more dark, then you will gain more PRG

- The AS spot is definitely clipped.

  8197   Thu Feb 28 03:25:27 2013 yutaUpdateLSCPR gain ~ 25 from PRMI carrier lock

[Manasa, Yuta]

We locked PRMI in carrier. Measured power recycling gain was ~25.

Plot:

  Here's some plot of PRC intra-cavity powers and MICH,PRCL error signals. As you can see from POPDC, cavity buildup was about 400, which means power recycling gain was ~25. Power recyling gain is fluctuating up to ~45 during lock. We need some gain normalization or something.
PRMIcarrier.png


Movie:

  Here's 30 sec movie of AS, POP, REFL when acquiring PRMI carrier lock. Although there's oscillation when acquiring lock, beam spot motion is less and stable compared with the past(before flipping PR2).



Locking details:
 == PRMI carrier ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg,  MICH_GAIN = -0.1, feedback to ITMX(-1),ITMY(+1)
  PRCL: REFL55_I_ERR, REFL55_PHASE_R = 70 deg, PRCL_GAIN = 5, feedback to PRM


Next:
  - Better filters and gains for stable lock
  - Kakeru method to measure g-factor (see elog around #1434)
  - OSA to measure g-factor

  8196   Thu Feb 28 02:43:49 2013 yutaUpdateLSCsome qualitative evidence of PRMI sideband lock

[Manasa, Yuta]

Since we have setup POP22 PD now(elog #8192), we could confirm that sideband power builds up when PRMI is sideband locked.

Plot:
  Here's some plot of PRC intra-cavity powers and MICH,PRCL error signals. As you can see from POP22, we locked at the peak of 11MHz sideband. There was oscillation at ~500 Hz, but we couldn't optimize the gain yet.
PRMIsideband.png


Movie:
  Here's 30 sec movie of AS, POP, REFL when acquiring (and losing) PRMI sideband lock. It was pretty hard to take a movie because it locks pretty seldom (~1 lock / 10 min).



Locking details:
  For MICH lock, we used ITMs instead of BS for reducing coupling between PRCL.
  Also, AS55 phase rotation angle was coarsely optimized by minimizing MICH signal in I.
  For PRCL lock, we used REFL55_I_ERR instead of REFL33_I_ERR. It had better PDH signal and we coarsely optimized phase rotation angle by minimizing PRCL PDH signal in Q.

 == PRMI sideband ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg,  MICH_GAIN = -0.1, feedback to ITMX(-1),ITMY(+1)
  PRCL: REFL55_I_ERR, REFL55_PHASE_R = 70 deg, PRCL_GAIN = -15, feedback to PRM

  We set POP22_PHASE_R = -170 deg by minimizing Q.

Issues:
 - We tried to use REFL55_Q_ERR to lock MICH, but couldn't. It looks like REFL error signals are bad.
 - We tried to use POP22_I_ERR to trigger PRCL lock, but it didn't work.

  8195   Wed Feb 27 23:19:54 2013 ranaUpdateSummary PagesMultiprocessing Implementation

  At first I thought that this was goofy, but then I logged in and saw that Megatron only has 8GB of RAM. I guess that used to be cool in the old days, but now is tiny (my laptop has 8 GB of RAM). I'll see if someone around has some free RAM for a 4600; in the meantime, I've killed a MEDM that was running on there and using up a few hundred MB.

Run your ssh-MEDMs elsewhere or else I'll make a cronjob to kill them periodically.

ELOG V3.1.3-